OpenStudy (pulsified333):

about independent events E and F in a sample space S, but assume that Pr[E]=0.2 and Pr[F]=0.4. Compute the following conditional probabilities: (1) Pr[E′|F]= (2) Pr[E|F′]=

OpenStudy (misty1212):

if they are independent, then multiply

OpenStudy (pulsified333):

hello again :D

OpenStudy (misty1212):

hi!!

OpenStudy (pulsified333):

but what happens if its the complement of one of them

OpenStudy (misty1212):

if they are independent then \[P(E'|F)=P(E')P(F)\]

OpenStudy (pulsified333):

oh sweet :D

OpenStudy (misty1212):

should be pretty clear that if the events are independent, then so are the compliments etc q

OpenStudy (pulsified333):

that didn't work

OpenStudy (pulsified333):

@misty1212

OpenStudy (misty1212):

hmm i see my mistake

OpenStudy (pulsified333):

oh

OpenStudy (misty1212):

to find\[P(E'|F)=\frac{P(E'\cap F)}{P(F)}=\frac{P(E')P(F)}{P(F)}=P(E')\] which is a silly way of saying that if E' and F are independent then \(P(E'|F)=P(E')\)